According to the outcome of the yesterday daily meeting, this morning I started to change the DAS powers both for NI and WI to match the values (see VIR-0698A-22) reported in the following table:
|DAS powers [mW]|
|WI IN||WI OUT||NI IN||NI OUT|
I encountered a problem with the WI DAS IN wave plate (LP2 in figure 1) as the corresponding calibrated channel (V1:TCS_WI_DAS_POWER_IN_CALI) did not reach values below 450 mW. Around 7.45 UTC, Suzanne and I went to the WI CO2 bench to verify that the power meter was working properly and whether the DAS calibration was still valid or not. After verifying that the DAS IN power meter was working well, we moved the Pickoff PM in front of the viewport and we switched on only the WI DAS IN to see if the value on the Pickoff corresponded to what reported by the calibrated channel: it was ok. Then we decided to check all the WI actuators and the NI actuators too, to be sure that the calibration was still valid for all the channels. All these checks went well. In the meanwhile we noticed that the V1:TCS_WI_DAS_POWER_IN_CALI decreased to 100 mW (due to the temperature decrease of the WI bench, as it was open - see figure 2).
When we came back to the control room, I tried to move again the WI DAS IN wave plate but the power could only increase, it would not go below around 130 mW.
We first investigated the possibility of a problem with the process TCSAgilisRot (even if it had low probability, since the other rotators were moving without problem). I stopped the process but I couldn’t get it to restart. I asked Nicola to try (see the operator entry 56364) and, at the end, the problem was fixed by Rhys Poulton and Franco Carbognani.
However, this did not solve the problem of the WI DAS IN, which, in the meanwhile, had increased to 450 mW, due to the bench thermalization after closing.
The other possibility is that 450 mW is really the minimum reachable power, even if on June 28th the power was set to 300 mW (as reported in logentry 56314). If this is the case, the solution is to move LP1 (see figure 1) which changes the power ratio between the two DAS arms. As a consequence, the WI DASes would need to be re-calibrated.
As decided at the daily meeting, tomorrow morning we will proceed with this option.